OpenStudy (anonymous):

Last year, McBarkie's Burgers contained 200 grams of fat in each burger. This year, the amount of fat has been reduced to 120 grams per burger. What is the percent of decrease in the number of grams of fat per burger to the nearest percent? A. 21% B. 2% C. 40% D. 20%

OpenStudy (anonymous):

Perecnt of decrease = change in # of grams/ original # of grams Perecent of decrease = 80/200 = 40/100 = 40% Choice (C) is correct.